# Isolated zero mode in a quantum computer from a duality twist

### v4

@@ III Detection of the unpaired zero mode in the IBM quantum device
In Fig. 2, we present both noiseless simulation and noisy quantum-hardware results for the bulk and defect autocorrelation functions obtained from the partial trace with 20 random basis states. We also show noiseless simulation results from Qiskit Aer simulators for 12-qubit and 20-qubit systems. The simulation results exhibit very small system-size dependence, indicating that the 20-qubit system accurately captures the thermodynamic limit. Furthermore, the results from the ibmq_kolkata device for the 20-qubit system is in good agreement with the noiseless simulations. We observe that the autocorrelation function of $\mathcal{O}=\sigma^{y}_{r}$ at the twist defect $r=L-1$ does not decay, which overlaps with the emergent Majorana zero mode. In contrast, all other autocorrelators decay to zero in the long time limit. The persistence of the autocorrelation function of the operator corresponding to the unpaired zero mode, during the Floquet dynamics, provides evidence for the dynamical generation of this mode in a digitized quantum device.
IV Unitary translation of twist defect
@@
A local unitary transformation of the form $U_{r}=CZ_{r-1,r}H_{r-1}$, where $CZ_{r-1,r}$ is a controlled-Z gate, and $H_{r-1}$ is a Hadamard gate, moves the twist-defect from the sites $(r-1,r)$ to sites $(r-2,r-1)$ [7, 30]. Fig. 3 illustrates this for a 4-qubit system. The unitary $U_{r}$ introduces a transverse field at site $r$ of strength $J$ while eliminating the transverse field at site $r-1$. It also changes the terms $J\sigma_{r-2}^{z}\sigma_{r-1}^{z}$ and $J\sigma_{r-1}^{z}\sigma_{r}^{x}$ to $J\sigma_{r-2}^{z}\sigma_{r-1}^{x}$ and $g\sigma_{r-1}^{z}\sigma_{r}^{z}$. Thus a Kramers-Wannier duality transformation has been performed on the sites $r,r-1$ where the roles of $g,J$ have been interchanged. Remarkably, the Majorana zero mode stays localized at $r$, and now separates two regions of the chain that are related by the Kramers-Wannier duality transformation. Although the Majorana zero mode stays localized at site $r$, in the language of Pauli strings, it is a more nonlocal object. In particular, instead of having an overlap with $\sigma^{y}_{r}$, it now has an overlap with $\sigma^{x}_{r-1}\sigma^{y}_{r}$.
@@
V Effects of Interactions
Now we add perturbations to the model (1) through the term
$H_{xx}=J_{x}\sum_{j}\sigma_{j}^{x}\sigma_{j+1}^{x}.$ (12)
@@ V Effects of Interactions
Note that $H_{xx}$ not only breaks integrability, but it also does not commmute with the discrete symmetry $\Omega_{r}$ due to the $\sigma^{x}_{r\pm 1}\sigma^{x}_{r}$ terms. We repeat the same quantum simulation for the interacting model with the defect location at site $r=L-1$. We find that the $\sigma^{y}_{r}$-correlator at the defect remains constant, while all other correlators decay with time as shown in Fig. 5. We observe a similar behavior to the noninteracting case shown in Fig. 2 for both bulk and defect autocorrelation functions, within the studied number of Floquet cycles.
@@
Figure 5: Signature of the Majorana zero mode at the defect site for duality-twisted interacting Floquet-TFIM. Panel (a) shows the nondecaying autocorrelation for $\sigma_{y}$ at the defect site $r=L-1$, which overlaps with the zero mode. All other autocorrelations both at the twist defect $r=L-1$ [panel (b), (c)] and at a bulk site $j=-1+L/2$ [panel (d), (e), (f)], far away from the twist defect, decay with time. The chosen bulk qubit behaves qualitatively similarly to other bulk qubits. We present data from ibmq_kolkata device for $L=20$, the longest physical loop of qubits available in the device. There is good qualitative agreement between noiseless simulations and measurements on the noisy quantum device. The noiseless simulations were also performed for a smaller system of 12 qubits. The agreement between the two system sizes $L=12,20$ indicates that we capture the behavior in the thermodynamic limit for these times. At longer times, we expect a stronger $L$ dependence discussed in the main text. The label $q_{i}$ indicates measurement of the $i$th qubit. Date of demonstrations is 22nd April, 2023.
@@
We note, however, that the long time dynamics of the $\sigma^{y}_{r}$-correlator at the defect is not captured in this figure as we can run the Floquet dynamics in the device only for about 20 Floquet cycles before noise takes over. We show the theoretical behavior of the Majorana zero mode by performing exact diagonalization for long times in Fig. 6(a). After the initial decay to the first plataeu, the Majorana mode decays to a second plateau. The plateau heights decrease with increasing interactions $J_{x}$. We capture only the initial portions of this behavior in the quantum device [see Fig. 6(b) and (c)], i.,e the decay to the first plateau. While the first plateau height is not $L$ dependent (as seen in the simulations in Fig. 5), we expect that the second plateau height is $L$ dependent, where as $L\rightarrow\infty$, the plateau height decreases to zero. This is because, there is no exact zero mode when we break integrability, with the zero mode decaying via scattering with bulk excitations [51]. However, a finite size system blocks these scattering processes leading to a partial decay, and hence a plateau height which is inversely related to system size [30]. A second important observation is that while the interactions now couple the two $Z_{2}$ sectors, yet the zero modes in the two sectors do not efficiently couple to each other as is evident by the appearance of the second plateau, which persists at long times. This is again a finite system size effect.
@@
Figure 6: Stability of the Majorana zero mode to integrability- and symmetry-breaking perturbations in a finite system. The autocorrelation function of $\sigma^{y}_{r}$ at the defect $r=L-1$, for different values of the perturbation $J_{x}$, are presented. Panel (a) shows the numerical results for system size $L=12$ up to large Floquet time steps. After the initial transient corresponding to the appearance of a first plateau, the autocorrelation stabilizes at a second plateau. This is a finite size effect where the height of the second plateau vanishes as $L\to\infty$ [30]. While we are not able to access this second plateau in the long-time limit in the ibmq_kolkata device due to noise, panels (b) and (c) show results from the IBM simulator and the ibmq_kolkata device, respectively, for intermediate timescales. At these timescales, the formation of the first plateau is clearly visible. Date of demonstrations is 27th April, 2023.
VI Conclusions
@@
In this paper, we investigated the physics of the duality twist defect and the corresponding emergent zero mode in a periodically driven TFIM loop of 20 qubits on the IBM superconducting quantum device. The interplay of topological effects and symmetries can lead to novel phenomena. The zero mode generated by a duality twist in a TFIM chain commutes with the $Z_{2}$ charge of the model and thus, theoretically, has an infinite lifetime in finite systems [18] (in contrast to edge states with open boundary conditions that are vulnerable to exponential hybridization). Using efficient random sampling, we obtained an accurate approximation to various autocorrelation functions and provided evidence, on the quantum computer, for the emergence of the unpaired zero mode bound to the duality twist defect. Furthermore, we explored the effects of moving the duality twist defect with unitary gates, and the effect of integrability- and symmetry-breaking perturbations. This work opens a new avenue for physically creating and investigating topological defects in quantum systems, and exploring their ramifications on nonequilibrium dynamics, a topic poorly understood theoretically.
@@
Acknowledgements.
We thank Oles Shtanko for helpful comments and discussions. We acknowledge the use of IBM Quantum services. We also thank the Brookhaven National Laboratory for providing access to IBM devices. This work was supported by the US National Science Foundation under Grants NSF DMR-2018358 (AM) and DMR-1945395 (AR and SS), and in part by the US National Science Foundation under Grant No. NSF PHY-1748958 (AR and AM).
@@ Appendix A Convergence of the partial trace
Figure 7: Autocorrelation functions from computing the partial trace in a classical computer for different sample sizes of random initial states, for system size $L=12$. The left panels [panel (a), (b), (c)] show the autocorrelation functions at the site of twist defect $r=L-1$, while the right panels [panel (d), (e), (f)] are for a site far away from it, $j=-1+L/2$.
All expectation values are the same, and correspond to the $Y$ autocorrelation function at the defect site. Thus for this particular observable, the trace is independent of the number of basis states used.
Appendix B Error Mitigation
We employ a composite error mitigation strategy combining error suppression and mitigation techniques [53]. Noisy bitstrings resulting from each shot are first mitigated by correcting readout errors using the method of Ref. 54, which utilizes calibration circuits scaling linearly with the number of qubits. Finally, zero-noise extrapolation is applied by extrapolating measured probabilities to the zero-noise value [55, 56]. This is achieved by performing measurements with 8,000 shots at noise factors of 1 (applying the original circuit) and 3 (evolving the system forward by the original circuit, backward by appending the inverse of the original circuit, and finally forward again by reapplying the original circuit [57]). A line through 1 and 3 on the horizontal axis yield the zero-noise extrapolated value at the intercept with the vertical axis.
Appendix C Layout of the computing device and its characteristics

### v6

@@ 3 Detection of the unpaired zero mode in the IBM quantum device
In Fig. 2, we present both noiseless simulation and noisy quantum-hardware results for the bulk and defect autocorrelation functions obtained from the partial trace with 20 random basis states. We also show noiseless simulation results from Qiskit Aer simulators for 12-qubit and 20-qubit systems. The simulation results exhibit very small system-size dependence, indicating that the 20-qubit system accurately captures the thermodynamic limit. Furthermore, the results from the ibmq_kolkata device for the 20-qubit system is in good agreement with the noiseless simulations. We observe that the autocorrelation function of $\mathcal{O}=\sigma^{y}_{r}$ at the twist defect r=L1r=L-1 does not decay, which overlaps with the emergent Majorana zero mode. In contrast, all other autocorrelators decay to zero in the long time limit. The persistence of the autocorrelation function of the operator corresponding to the unpaired zero mode, during the Floquet dynamics, provides evidence for the dynamical generation of this mode in a digitized quantum device.
4 Unitary translation of twist defect
@@
A local unitary transformation of the form $U_{r}=CZ_{r-1,r}H_{r-1}$, where $CZ_{r-1,r}$ is a controlled-Z gate, and $H_{r-1}$ is a Hadamard gate, moves the twist-defect from the sites (r1,r)(r-1,r) to sites (r2,r1)(r-2,r-1) [7, 32]. Fig. 3 illustrates this for a 4-qubit system. The unitary $U_{r}$ introduces a transverse field at site $r$ of strength $J$ while eliminating the transverse field at site r1r-1. It also changes the terms $J\sigma_{r-2}^{z}\sigma_{r-1}^{z}$ and $J\sigma_{r-1}^{z}\sigma_{r}^{x}$ to $J\sigma_{r-2}^{z}\sigma_{r-1}^{x}$ and $g\sigma_{r-1}^{z}\sigma_{r}^{z}$. Thus a Kramers-Wannier duality transformation has been performed on the sites r,r1r,r-1 where the roles of $g,J$ have been interchanged. Remarkably, the Majorana zero mode stays localized at $r$, and now separates two regions of the chain that are related by the Kramers-Wannier duality transformation. Although the Majorana zero mode stays localized at site $r$, in the language of Pauli strings, it is a more nonlocal object. In particular, instead of having an overlap with $\sigma^{y}_{r}$, it now has an overlap with $\sigma^{x}_{r-1}\sigma^{y}_{r}$.
@@
5 Effects of Interactions
Now we add perturbations to the model (1) through the term (other examples are presented in Appendix B)
$H_{xx}=J_{xx}\sum_{j}\sigma_{j}^{x}\sigma_{j+1}^{x}.$ (11)
@@ 5 Effects of Interactions
Note that $H_{xx}$ not only breaks integrability, but it also does not commmute with the discrete symmetry $\Omega_{r}$ due to the $\sigma^{x}_{r\pm 1}\sigma^{x}_{r}$ terms. We repeat the same quantum simulation for the interacting model with the defect location at site r=L1r=L-1. We find that the $\sigma^{y}_{r}$-correlator at the defect remains constant, while all other correlators decay with time as shown in Fig. 5. We observe a similar behavior to the noninteracting case shown in Fig. 2 for both bulk and defect autocorrelation functions, within the studied number of Floquet cycles.
@@
Figure 5: Signature of the Majorana zero mode at the defect site for duality-twisted interacting Floquet-TFIM. Panel (a) shows the nondecaying autocorrelation for $\sigma^{y}_{r}$ at the defect site r=L1r=L-1, which overlaps with the zero mode. All other autocorrelations both at the twist defect r=L1r=L-1 [panel (b), (c)] and at a bulk site j=1+L/2j=-1+L/2 [panel (d), (e), (f)], far away from the twist defect, decay with time. The chosen bulk qubit behaves qualitatively similarly to other bulk qubits. We present data from ibmq_kolkata device for $L=20$, the longest physical loop of qubits available in the device. There is good qualitative agreement between noiseless simulations and measurements on the noisy quantum device. The noiseless simulations were also performed for a smaller system of 12 qubits. The agreement between the two system sizes $L=12,20$ indicates that we capture the behavior in the thermodynamic limit for these times. At longer times, we expect a stronger $L$ dependence discussed in the main text. The label $q_{i}$ indicates measurement of the $i$th qubit. Date of demonstrations is 22nd April, 2023.
@@
We note, however, that the long time dynamics of the $\sigma^{y}_{r}$-correlator at the defect is not captured in this figure as we can run the Floquet dynamics in the device only for about 20 Floquet cycles before noise takes over. We show the theoretical behavior of the Majorana zero mode by performing exact diagonalization for long times in Fig. 6(a). After the initial decay to the first plateau, the Majorana mode decays to a second plateau. The plateau heights, especially of the second plateau, decrease with increasing interactions $J_{xx}$. We capture only the initial portions of this behavior in the quantum device [see Fig. 6(b) and (c)], i.e., the decay to the first plateau. While the first plateau height is not strongly $L$ dependent (as seen in the simulations in Fig. 5), we expect that the second plateau height is $L$ dependent, where as $L\rightarrow\infty$, the plateau height decreases to zero. This is because, there is no exact zero mode when we break integrability, with the zero mode decaying via scattering with bulk excitations [52]. However, a finite size system blocks these scattering processes leading to a partial decay, and hence a plateau height which is inversely related to system size [32]. A second important observation is that while the interactions now couple the two $\mathbb{Z}_{2}$ sectors, yet the zero modes in the two sectors do not efficiently couple to each other as is evident by the appearance of the second plateau, which persists at long times. This is again a finite system size effect. Here, we introduced a specific perturbation to the model that breaks its integrability and leads to a finite decay time for the Majorana mode. In Appendix B, we explore a broader class of perturbations that display similar behavior.
@@
Figure 6: Stability of the Majorana zero mode to integrability- and symmetry-breaking perturbations in a finite system. The autocorrelation function of $\sigma^{y}_{r}$ at the defect r=L1r=L-1, for different values of the perturbation $J_{xx}$, are presented. Panel (a) shows the numerical results for system size $L=12$ up to large Floquet time steps. After the initial transient corresponding to the appearance of a first plateau, the autocorrelation stabilizes at a second plateau. This is a finite size effect where the height of the second plateau vanishes as $L\to\infty$ [32]. While we are not able to access this second plateau in the long-time limit in the ibmq_kolkata device due to noise, panels (b) and (c) show results from the IBM simulator and the ibmq_kolkata device, respectively, for intermediate timescales. At these timescales, the formation of the first plateau is clearly visible. Date of demonstrations is 27th April, 2023.
6 Conclusions
@@
In this paper, we investigated the physics of the duality twist defect and the corresponding emergent zero mode in a periodically driven TFIM loop of 20 qubits on the IBM superconducting quantum device. The interplay of topological effects and symmetries can lead to novel phenomena. The zero mode generated by a duality twist in a TFIM chain commutes with the $\mathbb{Z}_{2}$ charge of the model and thus, theoretically, has an infinite lifetime in finite systems [18] (in contrast to edge states with open boundary conditions that are vulnerable to exponential hybridization). Using efficient random sampling, we obtained an accurate approximation to various autocorrelation functions and provided evidence, on the quantum computer, for the emergence of the unpaired zero mode bound to the duality twist defect. Furthermore, we explored the effects of moving the duality twist defect with unitary gates, and the effect of integrability- and symmetry-breaking perturbations. This work opens a new avenue for physically creating and investigating topological defects in quantum systems, and exploring their ramifications on nonequilibrium dynamics, a topic poorly understood theoretically.
@@
We thank Oles Shtanko for helpful comments and discussions. We acknowledge the use of IBM Quantum services. We also thank the Brookhaven National Laboratory for providing access to IBM devices. This work was supported by the US National Science Foundation under Grants NSF DMR-2018358 (AM) and DMR-1945395 (AR and $S$), and in part by the US National Science Foundation under Grant No. NSF PHY-1748958 (AR and AM).
@@ Appendix A Convergence of the partial trace
Figure 7: Autocorrelation functions from computing the partial trace in a classical computer for different sample sizes of random initial states, for system size $L=12$. The left panels [panel (a), (b), (c)] show the autocorrelation functions at the site of twist defect r=L1r=L-1, while the right panels [panel (d), (e), (f)] are for a site far away from it, j=1+L/2j=-1+L/2.
All expectation values are the same, and correspond to the $Y$ autocorrelation function at the defect site. Thus, for this particular observable, the trace is independent of the number of basis states used.
Appendix B Effects of other perturbations on stability
We show the stability of zero mode in the presence of various integrability- and symmetry-breaking perturbations. We restrict ourselves to only one-qubit and two-qubit terms for simplicity. We apply perturbations of the form $H_{\text{int}}=\lambda\sum_{j}h_{j}$, where $h_{j}$ is taken to be $\sigma^{y}_{j}$, $\sigma^{z}_{j}$, $\sigma^{x}_{j}\sigma^{y}_{j+1}$, $\sigma^{x}_{j}\sigma^{z}_{j+1}$, and $\sigma^{y}_{j}\sigma^{z}_{j+1}$. We note that $\sigma^{z}$ and $\sigma^{y}$ are nonlocal in terms of the Majorana operators as they contain Jordan-Wigner strings. The results for the $A_{Y}$ autocorrelation function are shown in Fig. 8.
Figure 8: Stability of the Majorana zero mode to various integrability- and symmetry-breaking perturbations in a finite size system of $L=12$, obtained from exact diagonalization. The autocorrelation function of $\sigma^{y}_{r}$ at the defect r=L1r=L-1 are presented as a function of Floquet steps, for different coupling constants. The autocorrelation functions decay to a non-zero value at long-times, where the plateau heights decrease with increasing coupling constants and also with increasing system sizes (not shown). In the thermodynamic limit $L\to\infty$, all the autocorrelations are expected to decay to zero.
Appendix C Error mitigation
We employ a composite error mitigation strategy combining error suppression and mitigation techniques [54]. Noisy bitstrings resulting from each shot are first mitigated by correcting readout errors using the method of Ref. [55], which utilizes calibration circuits scaling linearly with the number of qubits. Finally, zero-noise extrapolation is applied by extrapolating measured probabilities to the zero-noise value [56, 57]. This is achieved by performing measurements with 8,000 shots at noise factors of 1 (applying the original circuit) and 3 (evolving the system forward by the original circuit, backward by appending the inverse of the original circuit, and finally forward again by reapplying the original circuit [58]). A line through 1 and 3 on the horizontal axis yield the zero-noise extrapolated value at the intercept with the vertical axis.
Appendix D Layout of the computing device and its characteristics